# ScNeuGM: Scalable Neural Graph Modeling for Coloring-Based Contention and Interference Management in Wi-Fi 7

### v1

@@ II-B CSMA/CA with RTWT
We assume that all STAs and APs are synchronized in time, and the time is divided into slots indexed by $t$, where $t=1,2,\dots$. The duration of each slot is denoted as $\Delta_{\mathrm{0}}$ seconds. The network assigns a periodic slot $z_{k}$ to STA $k$ as its RTWT service period, where the same periodicity $Z$ is configured across all STAs in the network. The periodicity $Z$ directly affects the freshness of status updates from STAs. A higher value of $Z$ increases delays in status updates, whereas a lower value of $Z$ results in more STAs sharing the same slots, leading to contention and interference that can degrade transmission reliability. Specifically, STA $k$, for $k=1,\dots,K$, is scheduled to transmit exclusively during the periodic time slots $t=z_{k},z_{k}+Z,z_{k}+2Z,\dots$. Here, two STAs $i$ and $j$ transmit in the same slot if $z_{i}=z_{j}$, $\forall i,j$. For example, as shown in Fig. 3, three STAs ($K=3$) are assigned two slots within one RTWT period ($Z=2$). The illustrated case eliminates contention and interference between STAs $2$ and $1$ as well as between STAs $2$ and $3$, while contention and interference may occur between STAs $1$ and $3$.
In the allocated slot, each STA contends for channel access using CSMA/CA [9] when its assigned slots start periodically. In each assigned slot, the STA first checks whether any other STA is transmitting on the channel. If the channel is busy, the STA continues monitoring it until it becomes idle. Once the channel is idle, the STA initiates a random backoff period to further reduce the probability of collision with other STAs waiting to transmit. If the channel becomes busy during the backoff period, the countdown is paused and resumes once the channel is idle again. After the backoff period ends, the STA transmits one packet. If some STAs cannot sense this transmission (i.e., they are hidden from each other), simultaneous transmissions may occur, causing interference at the receiving AP and increasing the decoding error probability. The AP sends a negative acknowledgment if the packet transmission fails, triggering the STA to retransmit the same packet. Retransmission repeats until the maximum number of attempts is reached or an acknowledgment is received. Afterward, the STA waits for the next assigned slot in the subsequent period.
Due to contention from neighboring STAs, each STA may experience a significant backoff time, delaying the packet transmission to the later part of the slot. In this case, if transmitting the packet would occupy the channel beyond the ending time of the assigned slot, the STA skips the packet transmission in this slot and waits for the next assigned slot in the following period. Consequently, each STA contends and interferes only with those STAs that have the same assigned slots, while contention and interference are eliminated between STAs with different assigned slots.
@@ II-C Wi-Fi Network States
We next explain the network state information that will be used in the proposed slot assignment scheme.
@@ II-C1 Measured Network States
We assume that Wi-Fi devices have practical receiver sensitivities. In other words, if the path loss (in dB) between a STA and an AP exceeds a threshold $s_{\max}$ (in dB), the STA’s transmissions cannot be detected or received by the AP. Consequently, this path loss cannot be measured from the network. For a given $k$-th STA, where $k=1,\dots,K$, we collect and sort the APs that can detect the STA’s signal into a list as
@@
$\hat{\mathcal{S}}\triangleq\{\mathcal{S}_{1},\dots,\mathcal{S}_{K}\}.$ (6)
We refer to $\hat{\mathcal{S}}$ as the network states and $\mathcal{S}_{k}$ as STA $k$’s state, where $k=1,\dots,K$. Note that the size of each STA’s state, $\mathcal{S}_{k}$, may vary since STAs can have different numbers of APs capable of detecting their signals, depending on the locations of the STAs and APs. Here, $\hat{\mathcal{S}}$ represents the measured network states and serves as the input to the proposed methods.
@@ II-C2 Unmeasured Network States
Whether a pair of STAs is a contending or hidden pair significantly impacts their channel access and transmission reliability [11]. Contending and hidden STAs depend on the path losses between STA pairs, specifically on whether two STAs can detect each other. Measuring the path losses for all $K$ STAs requires approximately $\mathcal{O}(K^{2})$ measurements for all STA pairs, which introduces significant overhead to the network. Therefore, we assume that contending and hidden STA relationships are not measured during the online deployment of the proposed methods but are instead used as training information in offline simulations. Specifically, contending and hidden STA pairs are represented by binary indicators as
$O^{\mathrm{C}}_{i,j}\triangleq\mathbf{1}_{\{l(i,j)\leq\hat{l}\}};\ O^{\mathrm{H}}_{i,j}\triangleq\mathbf{1}_{\{l(i,j)>\hat{l}\}}\land\mathbf{1}_{\{s_{i,\hat{a}_{j}}\leq s_{\max}\}},\ \forall i,j.$ (7)
We set $O^{\mathrm{C}}_{i,j}=1$ if the distance $l(i,j)$ between STAs $i$ and $j$ is in the detectable range $\hat{l}$ of STA $j$, indicating STA $i$ is contending $j$, and set $O^{\mathrm{H}}_{i,j}=1$ if STA $i$ is not in the detectable range $\hat{l}$ of STA $j$ and STA $i$ interferes with STA $j$’s AP, i.e., STA $i$ is hidden from STA $j$. $O^{\mathrm{C}}_{i,j}+O^{\mathrm{H}}_{i,j}$ denotes that a STA pair $(i,j)$ is either contending or hidden STAs. We assume that if two STAs are neither contending nor hidden STAs, their transmissions do not affect each other.
III ScNeuGM Framework for Contention and Interference Management
This section begins by defining the QoS requirements of the STAs and formulating the RTWT slot assignment problem to manage contention and interference. Subsequently, the problem is transformed into a graph modeling task, where optimizing the edge values in the graph corresponds to optimizing its coloring scheme, which further optimizes the slot assignment. Finally, we present the ScNeuGM framework that trains the NN generating the optimal graph, highlighting the challenges posed in large networks.
III-A Contention and Interference Management Using RTWT
To improve QoS and reduce packet loss rates, we assign the slot choices of STAs $\mathbf{z}\triangleq[z_{1},\dots,z_{K}]$ in the period, where $z_{k}\in\{1,\dots,Z\}$ $\forall k$. Let $u_{k}(n)$ denote the number of packets successfully sent by STA $k$ during the $n$-th period, where $n=1,2,3,\dots$, and $u_{k}(n)\in\{0,1\}$ since each STA transmits only one packet per period, as mentioned in Section II. The reliability of each STA is defined as the successful packet transmission rate averaged over time as
@@ III-A Contention and Interference Management Using RTWT
Note that unsuccessful transmissions can occur due to insufficient transmission time in each assigned slot (e.g., caused by contention backoffs) or due to decoding errors (e.g., caused by interference). The QoS requirement is that each STA’s reliability must exceed a threshold $\hat{r}$. We express the expected value of the reliability of STA $k$ for given measured network states and slot assignment decisions as $\operatorname{\mathbb{E}}[r_{k}|\hat{\mathcal{S}},\mathbf{z}]$. The objective is to minimize the number of slots, $Z$, in the period or equivalently the transmission periodicity while ensuring transmission reliability as
$\min_{\mathbf{z},Z}\ Z,\ \text{s.t. }\operatorname{\mathbb{E}}[r_{k}|\hat{\mathcal{S}},\mathbf{z}]\geq\hat{r},\ z_{k}\in\{1,\dots,Z\},\ \forall k.$ (9)
The straightforward formulation in (9) is challenging to solve due to the lack of a clear structure for deriving the slot assignments and the minimum slot number from the network states. To address this issue, the problem is transformed into a graph modeling task, as described below.
III-B Problem Transformation to Graph Modeling Task
In Wi-Fi networks, contention and interference are pairwise interactions between STAs. To effectively represent these STA-pairwise interactions, we model the network as a binary-directed graph, $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where the STAs are vertices $\mathcal{V}$ and each directed edge $(i,j)\in\mathcal{E}$ implies that STA $i$’s transmissions will cause a QoS violation for STA $j$’s transmissions if they are in the same slot. The adjacency matrix $\mathbf{E}$ of $\mathcal{G}$ is defined with binary edges as
@@ III-B Problem Transformation to Graph Modeling Task
Graph coloring is employed to assign colors to vertices (STAs) such that no two adjacent vertices share the same color. Each color represents a distinct group of STAs that can transmit simultaneously without causing QoS violations, i.e., they can be assigned to the same slot. If an edge $(i,j)$ exists (i.e., $E_{i,j}=1$), STAs $i$ and $j$ must be assigned different colors to avoid contention and interference. We denote the coloring scheme of the binary graph with adjacency matrix $\mathbf{E}$ as
$\mathbf{z}^{\prime},Z^{\prime}=\chi(\mathbf{E}).$ (11)
Here, $\mathbf{z}^{\prime}=[z^{\prime}_{1},\dots,z^{\prime}_{K}]^{\rm T}$ represents the proper color choices for STAs $1,\dots,K$ that satisfy the edge constraints on the graph, and $Z^{\prime}$ denotes the minimum number of colors required to color the graph (i.e., the graph’s chromatic number), where $z^{\prime}_{k}\in\{1,\dots,Z^{\prime}\}$ $\forall k$. The coloring scheme is mapped to the slot assignments as follows.
@@
There exists an adjacency matrix $\mathbf{E}^{*}$ of the graph whose coloring scheme, $\mathbf{z}^{*},Z^{*}=\chi(\mathbf{E}^{*})$, solves the contention and interference management problem in (9). In other words, $\mathbf{z}^{*}$ and $Z^{*}$ represent the optimal slot assignments in (9), minimizing the number of slots while satisfying the STAs’ reliability constraints.
@@
The proof is in the appendix. ∎
Therefore, by optimizing the graph’s binary edge values and then coloring the graph, we can equivalently optimize the slot assignments in the problem (9). Based on this fact, we construct the following graph modeling optimization task as
[…]
$\mathbf{z},Z=\chi(\mathbf{E}),$
where the graph coloring in (11) is embedded to decide the STAs’ slots and the edge values of the graph are optimized instead. Since there exists an optimal graph whose coloring scheme is the optimal slot assignment, as shown in Proposition 1, we can find the optimal slot assignments in (9) by solving the problem (12) instead. As a result, this formulation in (12) transforms the slot assignment problem into a graph modeling task that finds the optimal binary graph representing the impact of the contention and interference on STAs’ transmissions.
III-C Neural Graph Modeling
Note that the optimal graph model is unknown and is difficult to derive analytically due to the complex interactions among STAs. Therefore, we simplify the problem by assuming that each binary edge value in the optimal graph, $E^{*}_{i,j}$, is a function of the states of STA $i$ and STA $j$, $\mathcal{S}_{i}$ and $\mathcal{S}_{j}$, as
@@ III-C Neural Graph Modeling
where we refer to $\mu^{*}(\cdot)$ as the optimal graph modeling function, mapping each STA pair’s states to the optimal edge value between them. This assumption further transforms the slot assignment task into the problem of finding the optimal graph modeling function $\mu^{*}(\cdot)$. Note that such a function can be heuristically designed based on human intuition, e.g., using the network state information as $\mu^{*}(\mathcal{S}_{i},\mathcal{S}_{j})\approx\mathbf{1}_{\{\mathcal{A}_{i}\cap\mathcal{A}_{j}\neq\emptyset\}},\forall i,j$ [21], where $\mathcal{A}_{i},\forall i$ can be computed using $\mathcal{S}_{i}$ as (3). However, heuristic designs rely on a fixed function and therefore cannot be adjusted to optimally capture the impact of contention and interference on STAs’ reliability or minimize the number of slots. This issue can be addressed by approximating the unknown structure of $\mu^{*}(\cdot)$ using a NN [35], as
@@
where $\theta^{\mu}$ denotes $\mu(\cdot)$’s parameters that can be flexibly trained using machine learning (ML). We refer to $\mu(\cdot|\theta^{\mu})$ as the NGM NN for Wi-Fi contention and interference management.
Figure 4: Iterative training of the NGM NN $\mu(\cdot|\theta^{\mu})$ for the contention and interference management problem in the Wi-Fi network.
Fig. 4 illustrates the iterative training of the NGM NN for managing contention and interference. The process begins with the NGM NN, $\mu(\cdot|\theta^{\mu})$, computing the edge values $\mathbf{E}$ between STA pairs to generate the binary graph model $\mathcal{G}$. This graph is then colored to obtain the slot assignment decision, $\mathbf{z}$ and $Z$. The network evaluates the performance of the decision and computes the loss, $L(\theta^{\mu})$, on the NN parameters $\theta^{\mu}$. Finally, the NN parameters $\theta^{\mu}$ are updated based on the loss value using the ML algorithm, and the workflow repeats until the NN is optimized. Note that graph coloring is an NP-hard task. To perform graph coloring efficiently with low complexity, we employ a greedy coloring scheme [42] that prioritizes STAs with higher degrees. Meanwhile, designing coloring algorithms is beyond the scope of this work.
III-D Overview of ScNeuGM Framework
When scaling the NGM to a large network with a significant number of Wi-Fi devices, several issues arise due to the massive amount of STA pairs, negatively impacting the NGM performance. The ScNeuGM framework addresses them accordingly as follows.
III-D1 Implicit Loss on Edge Values
In the iteration of the NGM training, the NN parameters $\theta^{\mu}$ are updated in the direction of minimizing the loss $L(\theta^{\mu})$, i.e., the gradient of the loss w.r.t. the parameters, $-\nabla_{\theta^{\mu}}L(\theta^{\mu})$. Typically, this gradient is estimated using the chain rule [32, 35], which decomposes into the gradient of the loss w.r.t. each edge value and the gradient of the edge value w.r.t. the NN parameters, e.g.,

### v2

@@ II-B CSMA/CA with RTWT
We assume that all STAs and APs are synchronized in time, and the time is divided into slots indexed by $t\in\{0,1,\ldots\}$. The duration of each slot is denoted as $\Delta_{\mathrm{0}}$ seconds. For RTWT service, denote by $Z$, which is a positive integer, the periodicity for all STAs, while $z_{k}\in\{1,\ldots,Z\}$ represents the offset of STA $k$ its RTWT service. The periodicity $Z$ directly affects the freshness of status updates from STAs. A larger $Z$ increases the delay between updates, while a smaller $Z$ causes more STAs to share the same slots, leading to higher contention and interference, which can degrade transmission reliability. Here, we assume that the slot duration $\Delta_{\mathrm{0}}$ is fixed while $Z$ can be adjusted to increase or decrease the total RTWT period duration $Z\Delta_{\mathrm{0}}$ to balance the update freshness and contention/interference levels. Specifically, STA $k\in\{1,\ldots,K\}$ is scheduled to transmit exclusively during the periodic time slots $t=z_{k},z_{k}+Z,z_{k}+2Z,\dots$. Note that multiple STAs can transmit in the same time slot if their transmission offsets are identical. For example, as shown in Fig. 3, suppose there are three STAs ($K=3$) and that $Z=2$. In this illustrated example, contention and interference between STAs 2 and 1, as well as between STAs 2 and 3, are eliminated, while contention and interference may still occur between STAs 1 and 3. Each STA contends for channel access using CSMA/CA [11] during its periodically assigned slots. The AP sends a negative acknowledgment if the packet transmission fails, triggering the STA to retransmit the same packet. Retransmission repeats until the maximum number of attempts is reached, an acknowledgment is received or the assigned slot has expired. Afterward, the STA waits for the next assigned slot in the subsequent period. Consequently, contention and interference occur only among STAs with the same offset, while such conflicts are avoided between those with different slot assignments.
@@ II-C Wi-Fi Network States
This subsection describes the network state information that will be used to discuss the proposed slot assignment scheme.
@@ II-C1 Measured Network States
In this paper, Wi-Fi devices are assumed to operate with practical receiver sensitivities. Specifically, if the path loss between a STA and an AP exceeds a threshold $s_{\max}$, the AP is assumed to be unable to detect or receive transmissions from the STA, and such path loss values cannot be measured by the network. To this end, for each STA, the APs capable of detecting its signal are identified and sorted into a list based on the measured path losses as follows:
@@
[…]
which is referred to as the collection of network states, with $\mathcal{S}_{k}$ representing the state of STA $k$. Note that the size of each STA’s state, $\mathcal{S}_{k}$, may vary since STAs can have different numbers of APs capable of detecting their signals, depending on the locations of the STAs and APs. Here, $\hat{\mathcal{S}}$ represents the measured network states and is maintained at the edge controller as the input to the proposed slot assignment methods.
@@ II-C2 Unmeasured Network States
Whether a pair of STAs is a contending or hidden pair significantly impacts their channel access and transmission reliability [13]. Contending and hidden STAs depend on the path losses between STA pairs, specifically on whether two STAs can detect each other. Measuring the path losses for all $K$ STAs requires approximately $\mathcal{O}(K^{2})$ measurements for all STA pairs, which introduces significant overhead to the network. Therefore, we assume that contending and hidden STA relationships are not measured during the online deployment of the proposed methods but are instead used as training information in offline simulations. Specifically, contending and hidden STA pairs, $\forall i,j$, are represented by binary indicators as
$O^{\mathrm{C}}_{i,j}\triangleq\mathbf{1}_{\{g(i,j)\leq s_{\max}\}};O^{\mathrm{H}}_{i,j}\triangleq\mathbf{1}_{\{g(i,j)>s_{\max}\}}\land\mathbf{1}_{\{s_{i,\hat{a}_{j}}\leq s_{\max}\}}.$ (7)
We set $O^{\mathrm{C}}_{i,j}=1$ if the path loss $g(i,j)$ between STAs $i$ and $j$ is less than or equal to the threshold $s_{\max}$, indicating STA $i$ is contending with STA $j$, and set $O^{\mathrm{H}}_{i,j}=1$ if the path loss $g(i,j)$ between STAs $i$ and $j$ exceeds $s_{\max}$ and STA $i$’s path loss to STA $j$’s associated AP, $s_{i,\hat{a}_{j}}$, is less than or equal to $s_{\max}$, where STA $i$ interferes with STA $j$’s AP and is hidden from STA $j$. $O^{\mathrm{C}}_{i,j}+O^{\mathrm{H}}_{i,j}$ denotes that an STA pair $(i,j)$ is either contending or hidden STAs. We assume that if two STAs are neither contending nor hidden STAs, their transmissions do not affect each other.
III The Proposed IGL Framework
In this section, we first introduce the problem formulation for contention and interference management. We then transform the problem into an interference graph-constructing task, where optimizing the edge weights corresponds to finding an optimal coloring scheme, which in turn determines the slot assignment. Finally, we present the IGL framework, which trains a neural network to generate the optimal graph structure while addressing the challenges posed by large-scale networks.
III-A Problem Formulation
To improve QoS and reduce packet loss rates, we can assign the slot choices of STAs $\mathbf{z}\triangleq[z_{1},\dots,z_{K}]$ in the period, where $z_{k}\in\{1,\dots,Z\}$. Let $u_{k}(n)$ denote the number of packets successfully sent by STA $k$ during the $n$-th slot. Note that $u_{k}(n)\in\{0,1\}$ since each STA transmits only one packet per slot, as mentioned in Section II. Then, the reliability of each STA is defined as the successful packet transmission rate averaged over time as
@@ III-A Problem Formulation
Note that unsuccessful transmissions can occur due to insufficient transmission time in each assigned slot (e.g., caused by contention backoffs) or due to decoding errors (e.g., caused by interference).
Suppose that we have a QoS requirement such that each STA’s reliability must exceed a threshold $\hat{r}$. Specifically, the expected value of the reliability of STA $k$, given the measured network states and slot assignment decisions, $\mathbb{E}[r_{k}|\hat{\mathcal{S}},\mathbf{z}]$, must be greater than or equal to $\hat{r}$, which becomes the constraint. The objective is to minimize the number of slots, $Z$, in the period or, equivalently, the transmission periodicity. Minimizing $Z$ is essential for preserving update freshness. A smaller $Z$ enables more frequent status reporting, whereas a larger one increases the risk of stale information, potentially leading to suboptimal control actions and safety-critical failures. In IIoT applications that require real-time monitoring and control, optimizing $Z$ is therefore fundamental to maintaining timely data collection and delivery from STAs [4, 5, 6]. Finally, the optimization problem is given by
$\min_{\mathbf{z},Z}\ Z$ (9)
$\text{s.t. }\ \operatorname{\mathbb{E}}[r_{k}|\hat{\mathcal{S}},\mathbf{z}]\geq\hat{r},\ z_{k}\in\{1,\dots,Z\},\ \forall k.$
However, the straightforward formulation in (9) is difficult to solve due to the lack of a clear structure for deriving the slot assignments and the minimum number of slots from the network states. To address this challenge, the problem is reformulated as a graph-constructing task, which will be discussed in the following section.
III-B Problem Transformation to Graph-Construction Task
In Wi-Fi networks, contention and interference are pairwise interactions between STAs. To effectively represent these STA-pairwise interactions, we model the network as a binary-directed graph, $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where the STAs are vertices $\mathcal{V}$ and each directed edge $(i,j)\in\mathcal{E}$ implies that STA $i$’s transmissions will cause a QoS violation for STA $j$’s transmissions if they are in the same slot. Specifically, the QoS violations, i.e., packet losses, happen 1) when one STA contends the channel with the other STA as they are close to each other, causing a significant backoff time and delaying the packet transmission until no sufficient time remaining in the slot; 2) when one STA is hidden from the other STA and is close to the other STA’s receiving AP, causing interference at the receiving AP. The adjacency matrix $\mathbf{E}$ of $\mathcal{G}$ is defined with binary edges as
@@ III-B Problem Transformation to Graph-Construction Task
Graph coloring is employed to assign colors to vertices (STAs) such that no two adjacent vertices share the same color. Each color represents a distinct group of STAs that can transmit simultaneously without causing QoS violations, i.e., they can be assigned to the same slot. If an edge $(i,j)$ exists (i.e., $E_{i,j}=1$), STAs $i$ and $j$ must be assigned different colors to avoid contention and interference.
Remark 1.
The graph can be also be constructed as a weighted graph, where the edge weights represent the severity of contention and interference between STAs. In this case, this STA is more likely to be separated from its neighboring STAs when the sum of its edge weights is large (implying a significant aggregated interference or contention). We can achieve the above separation rules by using weighted graph coloring, where the sum of edge weights from neighbors with the same color is required to be less than an interference threshold. For simplicity, we assume that the edge weights can only take 0 or $1$ and the interference threshold is $1$, considering the binary interference effect on QoSs and ignoring the aggregated small interference case, while the proposed methods can be straightforwardly extended to weighted graphs [37], utilizing weighted graph coloring algorithms [45].
We denote the coloring scheme of the binary graph with adjacency matrix $\mathbf{E}$ as
$\{\mathbf{z},Z\}=\chi(\mathbf{E}),$ (11)
where $\mathbf{z}=[z_{1},\dots,z_{K}]^{\rm T}$ is decided as the proper color choices for STAs $1,\dots,K$ that satisfy the edge constraints on the graph, and $Z$ now represents the minimum number of colors required to color the graph (i.e., the graph’s chromatic number), where $z_{k}\in\{1,\dots,Z\}$ $\forall k$. The coloring scheme is mapped to the slot assignments as follows.
@@
There exists an adjacency matrix $\mathbf{E}^{*}$ of the graph whose coloring scheme, $\{\mathbf{z}^{*},Z^{*}\}=\chi(\mathbf{E}^{*})$, solves the contention and interference management problem in (9). In other words, $\mathbf{z}^{*}$ and $Z^{*}$ represent the optimal slot assignments in (9), minimizing the number of slots while satisfying the STAs’ reliability constraints.
@@
See Appendix A ∎
Consequently, by optimizing the graph’s binary edge values and then coloring the graph, we can equivalently optimize the slot assignments in the problem (9). Based on this fact, we construct the following graph optimization task as
$\min_{E_{i,j}\in\{0,1\},\forall i\neq j}\ Z$ (12)
$\ \textrm{s.t.}\ \operatorname{\mathbb{E}}[r_{k}|\hat{\mathcal{S}},\mathbf{z}]\geq\hat{r},\ \forall k,$
$\quad\quad\{\mathbf{z},Z\}=\chi(\mathbf{E}).$
where the graph coloring in (11) is embedded to decide the STAs’ slots, and the edge values of the graph are optimized instead. Since there exists an optimal graph as
$\{\mathbf{z}^{*},Z^{*}\}=\chi(\mathbf{E}^{*}),$ (13)
whose coloring scheme is the optimal slot assignment, as shown in Proposition 1, we can find the optimal slot assignments in (9) by solving the problem (12) instead. As a result, this formulation in (12) transforms the slot assignment problem into an interference graph-constructing task that finds the optimal binary graph representing the impact of the contention and interference on STAs’ transmissions.
III-C Interference Graph Learning
We simplify the problem by assuming that each binary edge value in the optimal graph, $E^{*}_{i,j}$, is a function of the states of STA $i$ and STA $j$, $\mathcal{S}_{i}$ and $\mathcal{S}_{j}$, as [^1] In practice, the interference relationship may be influenced by other STAs in the network, especially in dense deployments. Future work can explore more complex models that consider higher-order interactions among multiple STAs when determining edge values, e.g., by incorporating GNN architectures to generate the graph edges.
@@ III-C Interference Graph Learning
where we refer to $\mu^{*}(\cdot)$ as the optimal graph-constructing function, mapping each STA pair’s states to the optimal edge value between them. This assumption further transforms the slot assignment task into the problem of finding the optimal graph-constructing function $\mu^{*}(\cdot)$.
In practice, the optimal graph model is unknown and is difficult to derive analytically due to the complex interactions among STAs and their associated APs. Therefore, the edge values have been heuristically designed based on human intuition on how contention and interference impact the network. For example, Edges can be designed as $\mu^{*}(\mathcal{S}_{i},\mathcal{S}_{j})\approx\mathbf{1}_{\{\mathcal{A}_{i}\cap\mathcal{A}_{j}\neq\emptyset\}},\forall i,j$ [31], where $\mathcal{A}_{i},\forall i$ can be computed using $\mathcal{S}_{i}$ as (3). However, human intuition-based designs rely on a fixed function and therefore cannot be adjusted to optimally capture the impact of contention and interference on STAs’ reliability or minimize the number of slots. Specifically, the following factors will impact the interaction between two STAs $i$ and $j$ and differentiate how one STA’s transmission affects the other STA’s transmission:
  • $O^{\mathrm{C}}_{i,j}$ and $O^{\mathrm{H}}_{i,j}$: Indicators on the STA pair being contending or hidden, where packet losses are caused by backoffs at transmitting STAs or low SINRs at receiving APs, respectively. As these indicators are not measured, we predict them based on measured path losses.
These factors vary every STA pair, and how they are quantitatively related to the QoS violations is difficult to derive. Thus, it is hard to design the optimal graph-constructing function manually. This issue motivates us to use IGL that approximates the unknown structure of $\mu^{*}(\cdot)$ using an NN [37], as
@@
where $\theta^{\mu}$ denotes $\mu(\cdot)$’s parameters that can be flexibly trained using machine learning (ML). We refer to $\mu(\cdot|\theta^{\mu})$ as the IGL NN for Wi-Fi contention and interference management.
Figure 4: Iterative training of the IGL NN $\mu(\cdot|\theta^{\mu})$ for the contention and interference management problem in the Wi-Fi network.
Fig. 4 illustrates the iterative training of the IGL NN for managing contention and interference. The process begins with the IGL NN, $\mu(\cdot|\theta^{\mu})$, computing the edge values $\mathbf{E}$ between STA pairs to generate the binary graph model $\mathcal{G}$. This graph is then colored to obtain the slot assignment decision, $\mathbf{z}$ and $Z$. The network evaluates the performance of the decision and computes the loss, $L(\theta^{\mu})$, on the NN parameters $\theta^{\mu}$. Finally, the NN parameters $\theta^{\mu}$ are updated based on the loss value using the ML algorithm, and the workflow repeats until the NN is optimized. Note that graph coloring is an NP-hard task. To perform graph coloring efficiently with low complexity, we employ a greedy coloring scheme [45] that prioritizes STAs with higher degrees. Meanwhile, designing coloring algorithms is beyond the scope of this work.
III-D Challenges of Scaling IGL in Large Wi-Fi Networks
When scaling the IGL to a large network with a significant number of Wi-Fi devices, several issues arise due to the massive number of STA pairs, degrading IGL performance. The proposed IGL framework, described in the following sections, addresses these issues as follows.
III-D1 Implicit Credit Assignment on Edge Values
In the iteration of the IGL training, the NN parameters $\theta^{\mu}$ are updated in the direction of minimizing the loss $L(\theta^{\mu})$, i.e., the gradient of the loss w.r.t. the parameters, $\nabla_{\theta^{\mu}}L(\theta^{\mu})$. Typically, this gradient is estimated using the chain rule [34, 37], which decomposes into the gradient of the loss w.r.t. each edge value and the gradient of the edge value w.r.t. the NN parameters as